OpenStudy (anonymous):

Find the values of x and y for which the lines are parallel. a)x = 47, y = 79 b)x = 58, y = 57 c)x = 79, y = 49 d)x = 79, y = 47

OpenStudy (anonymous):



OpenStudy (anonymous):

well because the lines are parallel, x-5 = 74 and 74+58+(y+1) = 180 ----> 133 + y = 180 so just solve for x and y now

OpenStudy (anonymous):

74=x-5 x=74+5=79 74+58+y+1=180 133+y=180 y=180-133=47